# LakeVilla: Multi-Table Transactions for Lakehouses

### v1

@@ 3.1. LakeVilla [Recovery] (LV [R])
[Technique 2: Logical and physical SAGA instructions for ”UPDATE ¡Table ID¿ SET … WHERE key=¡key¿”.]Technique 2: Logical and Physical SAGA instructions for ”UPDATE ¡Table ID¿ SET … WHERE key=¡key¿”. The physical instructions explicitly command reading, adding, and removing a specific file. Logical operations describe the logic of an update by searching for a key and replacing the found file.
Technique #2: Sublog Structure. After introducing our marker system, we propose an in-flight transaction recovery system using our sublogs. This technique generates a structure of transaction logs for in-flight transactions utilized by Undos and Redos of a transaction. We designed our sublog structure and its sublogs from scratch because the underlying operations differ in file-based OTFs from database systems. Operations generate, swap out, and split files to create a complete version for the following snapshot. Delta Lake’s time-traveling (Armbrust et al., 2020) already defines basic redo information for OTFs in its logs. However, its definition is static, while a transaction must adapt in a conflict. Hence, we need a new logging type for redo operations in a concurrent LH environment. LV defines physical and logical sublogs to provide the necessary information for this technique. Figure 2 shows the generated LV sublogs in our physical and logical definition for the query ”UPDATE ¡Table ID¿ SET … WHERE key=¡key¿”.
Our Physical Sublogs describe the explicit changes made to the previous OTF table state, such as adding and removing files. We show an example of a physical sublog for the above query in Fig. 2(a). The first instruction in Fig. 2(a) reads the file with the specified key ¡key¿ for the selected table ¡Table ID¿, resulting in the physical path of the found data file ¡path_old¿ and the file’s contents. The respective compensation step repeats this operation, as the redo operation does not change the overall table state. For the next step, the client uploads the new data file under the newly generated path ¡path¿ and defines its removal for the compensation step. Lastly, the transaction removes the old data file at ¡path_old¿ and adds it for compensation.
Our Logical Sublogs encode the example update query as shown in Fig. 2(b). Its first sublog defines the search of the current file with the given ¡key¿ and stores the result in the client’s buffer. The respective compensation step logs the old value of the buffer “old_found” to restore the old buffer state. The second sublog references the current buffer value “found” and merges the referenced file with a new file at ¡path¿ to generate the updated version. For compensation, we replace the new file at ¡path¿ with the original file. A transaction must ensure that it executes a redo on all its logical sublogs before it commits.
Sublog: File-based Redos and Undos. An undo traverses the sublog structure from the newest to the oldest sublog, executing the compensation steps to restore previous states. For our example in Fig. 2, it effectively restores the transaction state before the update in both logging versions: For the physical instructions, it adds the old file at ¡path_old¿ and removes the updated file at ¡path¿ (compensation steps p+1&p+2). Furthermore, the undo mechanism can execute physical reads like (compensation step p). The logical version starts by replacing ¡path¿ with the path stored in its buffer (compensations step l+1). Internally, we implement replacements through the removal of ¡path¿ and the addition of the file at the buffered path (compared to physical instructions). Lastly, we must store the old buffer value ¡old_key¿ in the client’s buffer for the following undo instructions (compensations step l). In contrast to undo, redo operations traverse the sublog structure from the oldest to the newest file, executing the redo steps. This process effectively results in a client performing the same operations on the current table version.
LV [R]’s Interplay of physical and logical sublogs. The goal of our sublog technique (#2) is to be adaptive to changed environments due to concurrency. In LV [R], physical sublogs describe file operations already performed on the object store and are thus not adaptive anymore. Therefore, they have precedence over logical sublogs. LV [R] reserves physical sublogs for operations without preceding in-flight transactions. If we detect a preceding transaction by its markers, we allow switching from physical sublogs to logical sublogs. However, it is not allowed to switch back because this would destroy the precedence. Furthermore, logical sublogs try to greedily apply explicit changes as much as possible to improve performance. For example, our design sees blind appends to a table as non-conflicting. A client will generate the file and add a physical sublog, even in a logical structure. This mixture of techniques #1 and #2 enables higher throughput in the metadata layer and transforms our markers from resembling exclusive to shared locks.
LV [R]’s Conflict resolution and file-granularity. To determine whether to commit or abort, LV [R] uses the markers placed in the metadata layer of a table. The unique and distinct order among the atomic markers ensures an ordering among all operations within a table. For concurrent transactions, the conflict detection triggers as soon as a preceding transaction replaces its marker. The waiting client accesses the content of the new log and checks if the prior transaction changed the same files or keys using its logical file-based sublog. LV [R] solves such file or key conflicts by undoing the transaction to the first operation where a conflict was detected and triggering a redo on the new table state. However, logical conflicts like dependencies across transactions or tables might trigger more prolonged recovery procedures, potentially completely re-executing the transaction. If an error appears multiple times, we added a retry counter that interrupts recovery if necessary and aborts the transaction. As OTF logs and snapshots operate on file references, this combination gives us file granularity for all transactions using LV[R].
LV [R]’s Client failure recovery. Besides our conflict resolution during commit, our LV [R] guarantees recovery from other causes of failure. Some common issues for OTFs revolve around upload/download failure, client crashes, or similar problems. In case of such failures, they might leave the LH metadata layer in an inconsistent or invalid state. Further, adding or removing files for updates or using markers and other LV techniques amplifies this problem. Hence, LV [R]’s sublogs are also a safety mechanism that allows clients to undo all failed transaction changes even if they did not initiate them. To guarantee recovery in such situations, LV[R] requires that sublogs can only removed after clients replace their marker with a snapshot/log. Currently, LV uses timeouts and counters for failed transactions to detect crashed clients with active markers. If a client decides to free the markers of a potentially crashed client, it does not delete the sublog. This rule enables the crashed client to automatically claim a new marker and continue the transaction.
3.2. LakeVilla [Complex Transactions] (LV [CT])
Multi-Query and Multi-Table Transactions in Lakehouses. Building upon LV [R], we implement multi-query and multi-table transactions for LHs. Re-applying the marker idea only partially implements complex transactions as multi-table transactions break global dependencies due to the independent metadata layers for all OTF tables. The metadata layer of one table would not be able to recognize any dependencies between tables, thereby enabling inconsistencies, deadlocks, and partial aborts. We avoid such problems in LV [CT] by combining our markers (technique #1) with a conflict detection model (technique #3): We demand a transaction must announce its global presence and check table compatibility using its markers.
Figure 3. Technique #3: A marker-based conflict and deadlock detection model. This figure shows a conflict between tables 1 and 2 and how T1 recovers from it.
[…]
[Technique 3: A marker-based conflict and deadlock detection model. This figure shows a conflict between tables 1 and 2 and how T1 recovers from it.]Technique 3: A marker-based conflict and deadlock detection model. This figure shows a conflict between Tables 1 and 2. T1 detects a conflict by reading T2’s markers. It frees its marker and places a new one to resolve the conflict.
Technique #3: Marker-based Conflict and Deadlock Detection. While a table-level lock and redo and undo information can prevent conflicts and dependencies within a table (LV [R]), we still can not capture dependencies between tables. For multiple tables, we can encounter a deadlock or cause inconsistencies when placing more markers into these tables and waiting from preceding transactions (compare deadlocks for locks). We track dependencies by introducing our deadlock and conflict detection for our OTF markers (technique #1) and define a global conflict and deadlock detection model. Our goal is to track dependencies across tables without causing deadlocks among transactions.
Detection Strategy. This technique requires transactions to acquire markers in every table accessed. Markers placed in tables must contain information about the respective transaction’s dependencies and concurrently check for changes in detected markers of other transactions. We explain our strategy with the example in Fig. 3: Transactions T1 and T2 start working on different tables and place their markers in the respective metadata layer. These transactions can be considered unrelated at this point. Later, T2 accesses Table 1 and places its marker behind T1’s marker. When initializing the new marker and reading the table state, T2 recognizes T1’s marker. As this relation does not violate its observed dependencies, T2 doesn’t abort. It updates all its lists of Tables 1 and 2 by overwriting each marker. T2 depends now on T1. Afterward, T1 accesses table 2, places its marker in the metadata layer of table 2, and checks for predecessors. T1 detects T2’s marker and retrieves its data. It compares the opened tables with T2’s list and checks for potential deadlocks and conflicts. Because T2 also lists table 1 as accessed and T1 is not aware of the ordering T1 depends on T2, it concludes that a deadlock might occur and triggers a marker shift. Hence, T1 frees its old marker in Table 1 by overwriting it with an empty manifest/log and places a new marker ”Marker T1’” into the respective metadata layer. After confirming that the dependencies match (also possibly readjusting some changes in Table 1), T1 continues its operation. If T2 decides to access another table and finds a conflict contradicting its internal observations, it will reevaluate this observation. Here, it checks the origin of this observation (here: Table 1) if the situation changes. Depending on the result, it updates its internal observations or solves the conflict.
LV [CT]’s Integration For multi-query transactions on the same table, we revisit the marker idea (technique #1) and extend it with technique #3. Like before, LV holds a marker to lock a table until the entire multi-query transaction is processed and overwrites with a real manifest/log to unlock (compare locks (Bernstein et al., 1987)). LV [CT] exposes information by allowing an in-flight transaction to claim access to new tables with multiple markers and using marker-based conflict detection. If a client detects a deadlock, it resolves the issues with technique #3’s marker shift. Furthermore, when successful in all tables, a commit triggers all table-level commits required for a transaction by replacing all markers and commits.
LV [CT] concurrency and delay handling. Due to the overhead of concurrent marker checks, marker generation, and general network delays, LV [CT] can not guarantee that all markers are up to date at any time. Hence, our file-based deadlock detection model must consider such uncertainties. We introduce a best-effort strategy: the concurrent checks to dependent markers searches for changes in dependencies. However, a transaction might be unable to detect the complete dependency cycle due to client crashes or faults. If a transaction is stuck waiting for other transactions, it triggers a marker shift after a timeout. After a timeout triggers a marker shift, the transaction aborts if a second timeout occurs. This strategy ensures that transactions do not hold their markers indefinitely, even under uncertainty, blocking the entire system.
Proof Sketch: Conflict Resolvement We demonstrate that LV [CT], combining techniques #1 and #3, effectively exposes and resolves conflicts across tables. From LV[R], we know that markers represent a running transaction within a table. They are unique for each transaction. As an LV [CT] client places a distinctive marker in each accessed table and waits for preceding ones, deadlocks can only occur if concurrent clients form a circular wait scenario across tables. Hence, if a transaction is part of such a cycle, it atomically places a marker behind another transaction’s marker, which waits for a different marker in another table. LV [CT] exposes its dependency list within all its markers, informing waiting transactions. At the same time, a client checks all markers it is waiting for to construct a dependency graph. If some given list contains the transaction’s ID or the provided information forms a cycle, an LV [CT] client is part of the cycle and triggers a marker shift. In case of larger cycles that might not be visible in the provided data, we use a timeout while waiting for markers to guarantee marker shifts breaking such cycles. Hence, we created a defensive approach that triggers shifts and potentially redos over too much wait. Performing a marker shift (technique #3) effectively frees all markers within such a cycle and claims newer ones. Hence, the marker shift inverts the relation between the transaction performing the shift and all waiting transactions. Therefore, during their concurrent checks, waiting transactions will realize that the marker is freed and proceed, breaking the cycle.
3.3. LakeVilla [Isolation] (LV [I])
The need for a global Lakehouse Layer. We introduce LV [I], defining a global layer that manages commits and global dependencies (technique #4). This change guarantees globally consistent versions at all times and global isolation (iii). Furthermore, it expands steps from the original OTFs and MVCC, like version assignment and validation (Bernstein and Goodman, 1983). Our approach minimizes cost and avoids slow requests like LIST to the object store.
Figure 4. Technique #4: A global version log tracking all valid snapshots. T2 succeeds and updates snapshots of Tables 3–5.
[…]
[Technique 4: A global version log tracking all valid snapshots. T2 succeeds and updates the snapshots of Tables 3, 4, and 5.]Technique 4: A global version log tracking all valid snapshots. The conflicting transactions T1 and T2 both enter the file-based validation phase. While T1 fails and is considered incompatible, T2 succeeds and updates the snapshots of Tables 3, 4, and 5.
Technique #4: Global Versioning After introducing our techniques to reserve snapshots, recover, and detect global deadlocks, we propose a global layer for object stores that manages all table snapshot versions to track and isolate transactions globally. Our global version log summarizes all tables’ most recent valid versions. If a client accesses a table, it is bound to the versions received by the global version log and can only use table versions equal to or smaller than the specified version. The example in Fig. 4 shows a global version log that tracks four tables and their newest snapshot versions. T1 and T2 both start their execution and retrieve the version combination (Table 1, Ver. 5), (Table 2, Ver. 3), (Table 3, Ver. 2), and (Table 4, Ver. 6) from the global version log.
Version Validation on the object store. To check the transaction’s compatibility with the current global version, we propose a file-based validation phase for the commit in OTFs, with a similar purpose as MVCC (Bernstein and Goodman, 1983; Neumann et al., 2015). This validation of a transaction starts by checking and atomically copying its version changes to a shared key on the object store. If the transaction succeeds, it iterates over its transaction version log and compares the version numbers of the current global version log. However, table versions in LHs are different. We can distinguish between a table’s read version (snapshot version we got from the first access of a table), hidden version (all versions created after a client received its read version), write version (id of the created log/marker), and log version (table version during validation). Using those abstractions, we define conflict rules for our global version log. We consider a transaction as compatible if it does not fulfill any of the following rules:
  • R1
    read version ¡¡ hidden version (outdated data)
  • R2
    write version ¡= hidden version (log number already taken)
  • R3
    log version + written logs != write version - written logs (lost update)
In summary, all of our rules have different purposes: R1 ensures that a read query does not access data that is too outdated from the current state. Further, R2 ensures that no log version is assigned twice, handles concurrency, and commits during transaction execution. Lastly, R3 guarantees the isolation with other transactions by only committing its logs. If a transaction is compatible, it executes the table-level commit of each accessed table before atomically announcing the new snapshots. In LV [I]’s snapshot isolation mode, only R2 and R3 are enforced, making LV [I] only check write sets.
LV [I]’s Read sets. We define our version log by using the well-proven read-and-write set abstraction. However, in LHs, we deal with table and file versions that must follow rules R1 – R3. For read transactions to retrieve globally consistent versions, they read the global version log and skip the snapshot assignment from their original OTF protocol. This change effectively reduces the number of requests to the metadata layers of single tables and replaces expensive requests (e. g., LIST for Delta Lake (Armbrust et al., 2020)). LV [I]’s read sets contain all tables read by a transaction. Hence, the LV [CT] transaction does not update such table versions. During its execution, we implement checking for concurrent changes in each table and updating hidden versions. In the serializable protocol, reads must pass the validation phase. Especially R1 was designed to uphold a configurable distance between the read and hidden versions. However, LV [I] can be relaxed to gain write serializability by automatically letting each read-only transaction pass validation and only incorporating the hidden version if necessary.
LV [I]’s Write sets. An LV [CT] write transaction follows the proposed strategy of technique #4. It retrieves a globally consistent read version by accessing the global version log and executes all its operations independently. Similar to reads, writes check for the hidden version of a table. A transaction’s write set contains all tables where it effectively created at least one temporary log entry. The respective validation checks its changes and atomically updates the global version log. To guarantee consistency, the transactions targeted write version number must be larger than their current log version, and the difference must match the number of added log files. LV [I] defines two special cases for writes in its validation: Updates and blind appends. Updates have to be read and written on the same table. Hence, they will have the same table in their read-and-write set. LV [I] enforces R1 – R3 on those sets independently and concurrently, making them the most work-intensive operations. On the other hand, blind appends only insert new data and are not bound by any restrictions. Hence, their order of operations is irrelevant. LV [I] implements a special flag for such operations that allows the transaction to take the following valid log ID directly in case of a validation failure.
Constructing an LV [I] version log. When accessing an OTF environment with LV [I] for the first time, the global version log does not exist. Hence, we construct this layer on demand using a mixture of the original OTF and technique #4: If an LV [I] client recognizes that no global version log exists, it proceeds by individually deducing its version of each accessed table. When committing, the transaction atomically enters the validation phase, accesses the empty global version log, and generates all entries based on its accessed table version. Other transactions accessing tables not contained in the version log will perform a similar mixture of claiming the version of the respective table themselves and adding the table during their validation. LV [I] defines adding a previously unknown table as always compatible because it can not check R1 – R3. When executed for a while, this procedure generates the global version log without disrupting other clients and replaces the version assignment proposed by technique #4.
LV [I]’s compatibility to other levels and original protocols. When using LV [I] alongside other client configurations, we face the problem of them ignoring the global version log. We ensure the compatibility of LV [I] by defining different behaviors for read and write transactions: Reads can still operate as described for technique #4. Such transactions do not conflict with concurrent transactions as they do not advance a table’s version number. In contrast, write transactions must be aware of such changes as concurrent non-LV [I] writes advance the version number without notifying the version log. If an LV [I] transaction decides to write to a table, it will check if the next possible version is available. If it detects a snapshot/log/marker of the same version, an inconsistency to the global version log occurs. It proceeds by reading the previously unknown logs, setting its hidden version to the newly read ID, and writing its results using the hidden version. During its commit, it counts the unknown log entries as it changes and commits them to the global version log if its initially assigned read version still equals the global version log. Otherwise, another concurrent transaction has already been committed using this version, and we must abort. Finally, to make writes using LV [I] visible to all clients, LV [I] still generates a valid log entry for each table.
4. LakeVilla’s Guarantees
After introducing the design of all LV features, we investigate their ACID guarantees summarized in Tab. 2. This section introduces the informal intuition behind individual proofs.
@@ 4. LakeVilla’s Guarantees
Table 2. Highest ACID guarantees achievable for OTFs + LV
@@ A OTF structure (ST) Atomic Markers (ST) Atomic Markers (ST) Version Validation
C Eventual (SQ,SW) Causal (ST) Causal Lineariz- ability
I Serializable (SQ,ST) (Databricks, 2024; Jain et al., 2023) Serializable (ST) Repeatable reads Serializable (LV)
@@ D Object store guarantee Object store guarantee Object store guarantee Object store guarantee
Single Query only: (SQ)  Single Table only: (ST)
Single Writer only: (SW)  LakeVilla only: (LV)
@@
Atomicity of LakeVilla. All LV Features guarantee table-level atomic and durable transactions as shown in Tab. 2. Atomicity revolves around the atomic markers in LV [R] and LV [CT]. Given the metadata structure of an OTF table, a transaction overwrites its marker with all changes made during its commit. Hence, any concurrent transaction will see either the ”empty” marker (see Sect. 3.1) or the complete log file. Because a marker is a single file, all changes of a transaction become visible at the same time when overwritten. However, this is limited to single tables due to the district metadata layers of OTF tables. In contrast, LV [I] guarantees Atomicity through its validation step, which uses atomic operations to allow only one client to modify the global version log. The head of the global version log is a single file that can be overwritten by a validated transaction in one operation, globally exposing all its changes simultaneously.
Durability of LakeVilla. Similar to the original OTF protocols, the guarantees of object stores and a client successfully uploading all new files before generating the log file (e. g., compare (Armbrust et al., 2020)) achieve Durability for LV [R] and LV [CT]. Further, a transaction only updates LV [I]’s global version log if it successfully enters validation and all tables’ metadata and data files are written. Hence, LV [I] uploads all data files to the object store before committing and can rely on the object store guarantees to provide Durability.
4.1. Consistency
@@
The Database Consistency of LakeVilla depends on the base OTF. Due to their metadata structure, Delta Lake and Iceberg offer ACID transactions for single tables (Projects, 2024b; Foundation, 2024a; Armbrust et al., 2020). All OTF transactions are ordered by the table’s ”chain” or ”tree” of the respective OTF metadata layer. Due to the atomic operations used (PUT-IF-ABSENT, Conditional Writes, extra service), each transaction claims a distinct spot in the metadata layer - defining the order of all successful transactions within the table. Because LV’s features keep this design and always lead to a valid state of the base OTF, all its features guarantee table-level consistency. Since OTFs are used in distributed settings and to clarify the feature’s consistency guarantees, we investigate the distributed consistency levels of each feature.
LV [R]: Causal Consistency (Single-Table). The causal consistency of LV [R] among clients using this feature is implied by its atomic markers. These markers reserve a snapshot version in the respective metadata layer and define an order between transactions. The operations of a transaction are already ordered. As clients writing with the original protocol might ignore markers, LV [R] checks for any logs written after its maker and aborts. However, LV [R] clients are not linearizable because write operations on different tables have no order among them.
@@
LV [CT]: Causal Consistency. LV [CT] provides causal consistency globally. A client’s operations are ordered by their execution order (related to the fact that they originate from the same client). Concurrent operations accessing the same table(s) are also related because they access the same data items (=tables). We define the order of such operations within a table using atomic markers. If a concurrent operation holds an earlier marker, all its operations are ordered beforehand; otherwise, all its operations are ordered after. However, LV [CT] is not serializable because it does not define an order among write operations on distinct table sets.
@@
LV [I]: Linearizability. Regarding LV [I]’s isolation level, we designed the validation process to be atomic. Hence, all write transactions must pass the validation sequentially, defining a global order. However, overlapping read-only transactions that retrieve their versions from the global version log during an active validation phase will work with outdated data. When they enter validation, they will likely succeed depending on the interval chosen for LV [I]’s R1. Hence, the read-only transaction must be ordered before the transaction processing the validation previously. This guarantees us a serial global history, not the order transactions committed. Hence, LV [I] guarantees Linearizability.
4.2. Isolation

### v2

@@ 5.1 LV[CT]’s Conflict Resolution
$\forall c\in C/{T_{x}},\forall s\in tr(c)\cap tr(T_{x}):w(s,c)<w(s,T_{x})$
Case 2: $T_{2}\rightarrow T_{1}$ does not close a cycle. Trivial, the resulting DAG is still acyclic, and $T_{2}$ can continue.
Proof Sketch: Conclusion. Hence, LV[CT] does not resolve in deadlocks during execution, given that the dependency DAG is complete. ∎
Handling Delays and Overhead in Practice. Due to the latency of object stores and to save costs, a client might not have the entire information available to build the entire DAG. Hence, it might be possible that larger cycles exist that remain undetected. For such cases, we implemented a timeout strategy for markers: A marker is guaranteed to only exist for a specified time. Further, if a transaction waits too long for another transaction, it preemptively triggers a marker shift, breaking potential larger cycles. In our LH use case, this mixture proved effective where LV[CT]’s module #3 resolves smaller cycles between transactions. In the rare case of concurrent, long-running transactions, marker timeouts prevent deadlocks and limit the DAG building overhead.
5.2 ACID Guarantees
@@
Table 2: Highest ACID guarantees achievable for OTFs by enforcing specific LV features
@@ A OTF structure (ST) Atomic Markers (ST) Atomic Markers (ST) Version Validation
C Eventual (SQ,SW) Causal (ST) Causal Linearizability
I Serializable (SQ,ST) isolation_databricks ; conf/cidr/0001KPDSZ23 Serializable (ST) Repeatable reads Serializable
@@ D Object store guarantee Object store guarantee Object store guarantee Object store guarantee
Single Query only: (SQ)    Single Table only: (ST)    Single Writer only: (SW)
LV allows transactions to choose specific features by temporarily implementing the requested LV feature. However, architects can build LHs with different ACID properties by enforcing different features on all transactions. We now explore these possibilities and prove their effectiveness. Our findings are summarized in Table 2.
5.2.1 Atomicity of LakeVilla
All LV Features guarantee table-level atomic transactions as shown in Table 2. Like the base OTFs, their Atomicity relies on atomic object store operations (if available), concurrent checks (see journals/pvldb/ArmbrustDPXZ0YM20 ), or strong consistency with PUT-IF-ABSENT operations aws_conditional_writes .
@@ 5.2.1 Atomicity of LakeVilla
For LV[R] and LV[CT], Atomicity revolves around their usage of atomic markers. Given the metadata structure of an OTF table, a transaction overwrites its marker with all changes made during its commit. Hence, any concurrent transaction will see either the marker (see Section 4.1) or the complete log file. Because a marker is a single file and each marker reserves a unique write snapshot for its transaction (see Equation 2), all changes of a transaction become visible at the same time when overwritten. However, this is limited to single tables due to the district metadata layers of OTF tables.
In contrast, LV[I] guarantees Atomicity through its validation step, which uses atomic operations to allow only one client to modify the global version log. The head of the global version log is a single file that can be overwritten by a validated transaction in one operation, globally exposing all its changes simultaneously.
5.2.2 Consistency of LakeVilla
This section discusses the database and distributed consistency levels of LV. It uses Definition 1 and extends it by:
Definition 2
We describe the ordering between two operations $op_{1},op_{2}$ using:
  • $op_{1}<op_{2}$: $op_{1}$ before $op_{2}$
  • $op_{1}>op_{2}$: $op_{1}$ after $op_{2}$
  • $op_{1}=op_{2}$: $op_{1}$ is concurrent to $op_{2}$
  • $op_{1}\sim op_{2}$: ${<,>,=}\sim\in\{$<,>,=\}
  • $op_{1}\not\sim op_{2}$ no order between $op_{1}$ and $op_{2}$
Equivalently, for two transactions $T_{1},T_{2}\in T$.
@@ 5.2.2 Consistency of LakeVilla
The Database Consistency of LV depends on the base OTF. Due to their metadata structure, Delta Lake and Iceberg offer ACID transactions for single tables journals/pvldb/ArmbrustDPXZ0YM20 ; Iceberg_website ; delta_doc . All OTF transactions are ordered by the table’s ”chain” or ”tree” of the respective OTF metadata layer. Due to the atomic operations used (PUT-IF-ABSENT, Conditional Writes, extra service), each transaction claims a distinct spot in the metadata layer and places all operations of a transaction in a single log file - defining the order of all successful transactions and their operations within a table $x$.
Lemma 2
OTFs guarantee an ordering of all transactions and their operations on the same table:
$\forall T_{1},T_{2}\in T:tr(T_{1})\cap tr(T_{2})\not=\emptyset\Leftrightarrow T_{1}\sim T_{2}\Leftrightarrow\forall op_{1}\in T_{1},op_{2}\in T_{2}:op_{1}\sim op_{2}$
Because LV’s features keep the OTF design and always lead to a valid state of the base OTF, all its features guarantee table-level consistency.
LV[R]: Causal Consistency (Single-Table). The causal consistency of LV[R] among clients using this feature is implied by its atomic markers and transactions, which always publish all their operations in a single file. Combined with the existing ordering mechanisms of OTFs (Lemma 2), they guarantee a logical and consistent ordering between transactions and their operations.
Proof Sketch: Hypothesis. LV[R] guarantees causal consistency for single-table, multi-query transactions.
Proof Sketch: Direct Proof Following Definition 1, all transactions $T_{i}\in T$ define an order among their operations. Hence:
$\forall T_{i}\in T,\forall x,y\in T_{i},x\not=y:x\sim y$ (3)
LV[R]’s marker technique guarantees that all $ops\in T_{i}\in T$ affecting the same table are placed in the same metadata file (Section 5.2.1). Following Lemma 2, the OTF layer defines an order between transactions on the same table (e.g., in Delta Lake: log files). Thus:
$\forall T_{1},T_{2}\in T,|tr(T_{1})|=1,tr(T_{1})=tr(T_{2}):T_{1}\sim T_{2}$ (4)
As LV[R] is applied to the base OTF, combining Equation 3 and Equation 4 gives:
Lemma 3
LV[R] guarantees causal consistency on a table level using its marker technique:
$\forall T_{1},T_{2}\in T,|tr(T_{1})|=1,tr(T_{1})=tr(T_{2}),x\in T_{1},y\in T_{2},x\not=y:x\sim y$
Proof Sketch: Conclusion. This implies table-level causal consistency for LV[R] as only the order of related operations is defined. This result was expected as LV[R] focuses on table-level guarantees and does not track dependencies between tables like the base OTFs ∎
@@
LV[CT]: Causal Consistency. LV[CT] provides causal consistency globally. A client’s operations are ordered by their execution order (Equation 3). Concurrent operations accessing the same table(s) are also related because they access the same data items ($|tr(T_{1})\cap tr(T_{2})|\not=0,$). We define the order of such operations within a table using atomic markers like LV[R] (compare Equation 4). Unlike LV[R], LV[CT]’s conflict detection builds a DAG of dependencies for the respective transaction, guaranteeing a consistent global ordering between such transactions and their operations, lifting the restriction to single-table transactions:
Lemma 4
LV[CT] guarantees causal consistency globally using its marker and conflict detection techniques:
$\forall T_{1},T_{2}\in T,|tr(T_{1})\cap tr(T_{2})|\not=0,x\in T_{1},y\in T_{2},x\not=y:x\sim y$
However, LV[CT] is not serializable because it does not define an order among operations on distinct table sets.
@@
LV[I]: Linearizability. Regarding LV[I]’s isolation, we designed the validation process to be atomic. Hence, all write transactions must pass the validation sequentially, defining a global order similarly to LV[R]’s marker on table level. However, overlapping read-only transactions that retrieve their versions from the global version log during an active validation phase will work with outdated data. Logically, when read-only transactions enter validation, they will likely succeed depending on the interval chosen for LV[I]’s R1. Hence, the read-only transaction must be ordered before the transaction that previously processed the validation. This guarantees a serial ordering between all transactions and, thus, Linearizability.
Lemma 5
LV[I] guarantees Linearizability globally using its global version log techniques:
$\forall T_{1},T_{2}\in T,x\in T_{1},y\in T_{2},x\not=y:x\sim y$
5.2.3 Isolation